# Multi-Agent Taskforce Collaboration: Self-Correction of Compounding Errors in Long-Form Literature Review Generation

### v2

@@ Abstract
Compounding error is critical in long-form literature review generation, where minor inaccuracies cascade and amplify across subsequent steps, severely compromising the faithfulness of the final output. To address this challenge, we propose the Multi-Agent Taskforce Collaboration (MATC) framework, which proactively mitigates errors by orchestrating LLM-based agents into three specialized taskforces: (1) an exploration taskforce that interleaves retrieval and outlining using a tree-based strategy to establish a grounded structure; (2) an exploitation taskforce that iteratively cycles between fact location and draft refinement to ensure evidential support; and (3) a feedback taskforce that leverages historical experience for self-correction before errors propagate. Experimental results show that MATC achieves state-of-the-art performance on existing benchmarks (AutoSurvey and SurveyEval), significantly outperforming strong baselines in both citation quality (e.g., +15.7% recall) and content quality. We further contribute TopSurvey, a new large-scale benchmark of 195 peer-reviewed survey topics, on which MATC maintains robust performance, demonstrating its generalizability.
Multi-Agent Taskforce Collaboration: Self-Correction of Compounding Errors in Long-Form Literature Review Generation
Zhi Zhang1, Yan Liu1thanks: Corresponding author., Zhejing Hu1, Gong Chen2, Sheng-hua Zhong3, Jiannong Cao1 1Department of Computing, The Hong Kong Polytechnic University, Hong Kong, China 2FireTorch Partners, Shenzhen, China 3College of Computer Science and Software Engineering, Shenzhen University, Shenzhen, China {zhi271.zhang, zhejing.hu}@connect.polyu.hk, {yan.liu, jiannong.cao}@polyu.edu.hk, heinz@clozzz.com, csshzhong@szu.edu.cn
@@ 1 Introduction
Early approaches primarily leverage multi-document summarization techniques to generate summaries from a set of reference papers Hoang and Kan (2010); Hu and Wan (2014); Erera et al. (2019), enabling the automatic creation of related work sections given a set of references as input. In recent years, large language models (LLMs) demonstrate impressive planning and reasoning abilities Wang et al. (2024a); Zhao et al. (2024); Zhong and Wang (2024); Nandy and Bandyopadhyay (2025); Zhang et al. (2025). This leads to a growing research area that employs LLMs as controllers to construct autonomous agents capable of effectively performing complex tasks Wang et al. (2024a). Building on these advances, recent state-of-the-art research, e.g., AutoSurvey Wang et al. (2024b) and SurveyX Liang et al. (2025), achieves remarkable success by designing LLM-based autonomous agents to automate the entire pipeline of literature review and generate long-form manuscripts.
Although these advances enable the automatic generation of human-like literature reviews with reduced manual effort, previous studies on multi-step models show that the inherent long-horizon workflow of such systems inevitably introduces the risk of compounding errors. In the workflow, small mistakes made at earlier steps can accumulate and propagate through subsequent steps, as errors that occur at a certain step will affect the next state Wang et al. (2025a). Without effective self-correction mechanisms, these deviations will further expand the error in later steps Han and Zhang (2023); Venkatraman et al. (2015); Hejna et al. (2023).
@@
This paper investigates the phenomenon of error accumulation in automatic literature review generation. As a demonstration, we employ the state-of-the-art large language model GPT-4.1 OpenAI (2025) to conduct a literature review generation workflow on the topic of EEG (electroencephalogram)-based emotion recognition, with a target length of 4,000 words. As shown in Fig. 1, when the retrieval stage fails to identify relevant literature, the resulting errors can mislead the outlining process. An irrelevant outline may further cause the drafted manuscript to deviate from the intended topic. In long-form text generation, the increased complexity and the introduction of more dependent steps further exacerbate the compounding errors.
To tackle this challenge, we propose the Multi-Agent Taskforce Collaboration (MATC) framework. MATC consists of two types of autonomous agents: the manager agent and the executor agents. The manager agent orchestrates the workflow, while the executor agents each specialize in a distinct step of the process, such as literature searching, outline generation, fact localization, or manuscript drafting.
Inspired by organizational quality management Hackman and Wageman (1995), we propose to orchestrate agents organized as taskforces. A taskforce refers to a temporary group of specialized agents focused on specific subtasks Altier (1987); Aikin (1949); de Waard and Kramer (2008). We formulate three types of taskforces, each designed with specific error mitigation mechanisms.
Given a user’s instruction, the manager agent oversees the entire workflow and initiates the formation of different taskforces. To mitigate error accumulation between searching and outlining agents, the manager organizes an exploration taskforce employing a tree-based strategy. The taskforce begins with a broad overview and incrementally determines the literature and outline at each level, preventing errors in ungrounded outlines or biased retrieval. To mitigate error accumulation between locating and drafting agents, the manager organizes an exploitation taskforce employing an iterative cycle: the draft guides fact location from the literature, while the located facts inform draft refinement, preventing errors in unsupported claims or mismatched evidence. To mitigate errors within each agent’s workflow, the manager organizes feedback taskforces that maintain historical experience and implement dynamic checklists to guide self-correction before errors propagate to subsequent stages. Finally, the manager integrates the outputs of all taskforces to produce the final manuscript.
@@ 2 Related Work
Automating the summarization of research papers has long attracted interest. Early work mainly used multi-document summarization techniques. In 2010, Hoang et al. introduced automatic related work summarization, extracting sentences from multiple articles to form an extractive summary tailored to the target paper Hoang and Kan (2010). Lu et al. constructed Multi-XScience, a large-scale dataset for generating related work sections based on a paper’s abstract and cited articles, serving as a benchmark for later studies Lu et al. (2020). Chen et al. proposed a relation-aware multi-document encoder and a relation graph to capture cross-document dependencies for abstractive related work generation Chen et al. (2021). Xiao et al. developed PRIMERA, a pre-trained encoder-decoder model with novel objectives for linking and aggregating information across documents Xiao et al. (2022). Liu et al. proposed aligning sentences by category and using sparse transformers to organize information from many references Liu et al. (2023).
Figure 2: The overall framework of Multi-agent Taskforce Collaboration (MATC) for literature review generation.
Recently, advances in LLMs have greatly improved the scope and quality of automated literature reviews, enabling end-to-end systems to generate literature review papers. Wang et al. introduced AutoSurvey, a framework with retrieval, outline generation, parallel subsection drafting, integration, and rigorous evaluation Wang et al. (2024b). Shao et al. developed STORM, which models the pre-writing stage by discovering diverse perspectives, simulating expert conversations, and curating information to create an outline used to generate a full Wikipedia article Shao et al. (2024). Liang et al. designed SurveyX, an end-to-end solution for automated survey generation, covering online literature search, organization, and survey writing Liang et al. (2025). While these methods represent impressive progress, little research has addressed the risk of errors in generated outputs. Self-correction algorithms in LLMs are not suitable for the workflow of literature review generation Madaan et al. (2023); Shinn et al. (2023); Ferraz et al. (2024). As systems become more complex, error accumulation becomes a significant challenge.
@@ 3 Methodology
As illustrated in Fig. 2, given a user instruction $U$, the proposed Multi-Agent Taskforce Collaboration (MATC) framework orchestrates three taskforces to produce the final literature review output $Z$. In the following sections, we introduce each taskforce in detail.
3.1 Exploration Taskforce
The exploration taskforce is designed to determine the outline of the literature review and to retrieve relevant references for each section of the outline. Without prior knowledge of the relevant literature, the initial draft outline tends to be suboptimal. Conversely, without a determined direction, the literature retrieval process lacks focus. To mitigate compounding errors, we propose a tree-based strategy.
In detail, given a user instruction $U$, the manager $A_{M}$ initiates the exploration taskforce. The manager constructs a tree with the user instruction as the root node, representing the overall direction for exploring the field. This tree is expanded by identifying sub-directions of interest. For the root node, $A_{M}$ first invokes the searching agent to retrieve literature based on the user instruction, resulting in a set of relevant papers. The searching agent $A_{S}$ operates as follows:
@@ 3.1 Exploration Taskforce
where $L_{i}^{(0)}$ denotes the $i$-th piece of literature retrieved for the root node at tree depth $d=0$. The number of literature is denoted by $I$.
Next, the manager assigns the outlining agent to determine sub-directions based on the titles and abstracts of the retrieved literature. The outlining agent $A_{O}$ operates as follows:
$\{O_{1}^{(0)},O_{2}^{(0)},\ldots,O_{J}^{(0)}\}=A_{O}(\{L^{(0)}_{i}\})$ (2)
where $O_{j}^{(0)}$ denotes the $j$-th sub-direction at the root level, and $J$ is the total number of sub-directions at tree depth $d=0$.
For each outline node $O_{j}^{(d)}$ at depth $d$, the manager decides whether further decomposition is needed according to the following criterion:
@@
$\{L^{(d+1)}_{1},L^{(d+1)}_{2},\dots,L^{(d+1)}_{I}\}$ $=A_{S}(O_{j}^{(d)})$ (4)
$\{O_{1}^{(d+1)},O_{2}^{(d+1)},\ldots,O_{\mathcal{J}}^{(d+1)}\}$ $=A_{O}(\{L_{i}^{(d+1)}\})$
@@
where $\{L_{i}^{(d+1)}\}$ denotes the set of literature retrieved for sub-direction $O_{j}^{(d)}$ at depth $d+1$, and $\mathcal{J}$ denotes the number of sub-directions.
@@
This process constructs child nodes in the tree $\mathcal{T}$, where nodes $\{O_{1}^{(d+1)},O_{2}^{(d+1)},\ldots,O_{\mathcal{J}}^{(d+1)}\}$ are the children of node $O_{j}^{(d)}$. The exploration process terminates when the following condition is met:
$\forall v\in\mathcal{V}:\phi(v)=0\quad\text{or}\quad d=d_{\max}$ (5)
@@
where $\mathcal{V}$ denotes all leaf nodes in the outline tree and $d_{\max}$ is the maximum allowed depth. The resulting tree $\mathcal{T}$ provides a hierarchical outline, with associated literature at each level, for downstream generation.
3.2 Exploitation Taskforce
@@
The exploitation taskforce is designed to extract relevant snippets from the literature and draft the review manuscript. Without clear target content, it is difficult to locate the necessary evidence in the literature. Conversely, without sufficient supporting evidence, the manuscript may lack grounding. To mitigate compounding errors, we propose a tree-based strategy.
@@
In detail, given the outline tree $\mathcal{T}$ produced as mentioned above, the manager $A_{M}$ initiates the exploitation taskforce. For each leaf node in the tree, $A_{M}$ identifies its parent outline node, denoted as $\{\mathcal{O}_{j}\}$, and the associated literature set $\{\mathcal{L}_{i}\}$. The exploitation taskforce then proceeds through iterative cycles. At the initial iteration $t=0$, the locating agent $A_{L}$ extracts relevant information based on the outline node $\{\mathcal{O}_{j}\}$ and the literature set $\{\mathcal{L}_{i}\}$, and operates as follows:
$\{F_{1}^{(0)},F_{2}^{(0)},\ldots,F_{N}^{(0)}\}=A_{L}(\{\mathcal{L}_{i}\},\{\mathcal{O}_{j}\})$ (6)
@@
where $F_{n}^{(0)}$ denotes the $n$-th note snippet extracted from the literature set $\{\mathcal{L}_{i}\}$ relevant to the parent outline $\{\mathcal{O}_{j}\}$, and $N$ is the total number of snippets.
@@
The drafting agent $A_{W}$ composes the initial draft as follows:
$M^{(0)}_{k}=A_{W}(\{F^{(0)}_{n}\},\{\mathcal{O}_{j}\})$ (7)
@@
where $M^{(0)}_{k}$ represents the initial manuscript for the outline set $\{\mathcal{O}_{j}\}$, incorporating the content of all its child outline nodes.
@@
For subsequent iterations $t>0$, the locating agent $A_{L}$ re-examines the literature to extract new snippets that support or improve the current draft, operating as follows:
$\{F_{1}^{(t)},F_{2}^{(t)},\ldots,F_{N}^{(t)}\}=A_{L}(\{\mathcal{L}_{i}\},M^{(t)}_{k})$ (8)
@@
where $M^{(t)}_{k}$ denotes the draft content from the previous iteration, and $\{F^{(t)}_{n}\}$ is the set of newly extracted note snippets. The drafting agent $A_{W}$ refines the section draft by incorporating the new evidence and the previous draft:
$M^{(t+1)}_{k}=A_{W}(\{F^{(t)}_{n}\},\{\mathcal{O}_{j}\},M^{(t)}_{k})$ (9)
where $M^{(t+1)}_{k}$ is the updated content, and $M^{(t)}_{k}$ serves as the previous context for refinement.
@@
The exploitation process for each parent node $\mathcal{O}_{i}$ terminates when the following condition is met:
$\psi(M^{(t+1)}_{i},M^{(t)}_{i})>\epsilon\quad\text{or}\quad t=t_{\max}$ (10)
@@
where $\psi(\cdot,\cdot)$ denotes the similarity between consecutive iterations of the draft, measured by the ROUGE-1 score Lin (2004), $\epsilon$ is a similarity threshold that indicates no further significant improvements can be made, and $t_{\max}$ is the maximum number of allowed iterations.
@@
The final content $M^{(t^{*})}_{i}$, where $t^{*}$ denotes the stopping iteration for parent node $i$, serves as the final draft. Finally, the manager $A_{M}$ aggregates the finalized drafts for all sections and delivers the completed literature review to the user.
3.3 Feedback Taskforce

### v3

# Structure-Guided Memory Consolidation for Mitigating Compounding Errors in Literature Review Generation

@@ Abstract
Compounding errors pose a significant challenge in automatic literature review generation, as inaccuracies can cascade across multi-stage retrieval and generation workflows. Existing self-correction strategies often lack mechanisms to effectively track and consolidate verified information throughout the process, making it difficult to prevent error accumulation and propagation. In this paper, we propose Structure-Guided Memory Consolidation (SGMC), a novel framework that incrementally consolidates and verifies information using structured representations at each stage of the literature review pipeline. SGMC consists of three key modules: Tree-Guided Memory for hierarchical literature retrieval and outline generation, Hub-Guided Memory for evidence extraction and iterative content refinement, and Self-Loop Memory for proactive error correction via historical feedback. Extensive experiments on public benchmarks and a newly constructed large-scale dataset demonstrate that SGMC achieves state-of-the-art performance in citation accuracy and content quality, significantly mitigating compounding errors in long-form literature review generation.
footnotetext: *Corresponding author
@@ 1 Introduction
Early approaches primarily leverage multi-document summarization techniques to generate summaries from a set of reference papers Hoang and Kan (2010); Hu and Wan (2014); Erera et al. (2019), enabling the automatic creation of related work sections given a set of references as input. In recent years, large language models (LLMs) have demonstrated impressive planning and reasoning abilities Wang et al. (2024a); Zhao et al. (2024); Zhong and Wang (2024); Nandy and Bandyopadhyay (2025); Zhang et al. (2025). This has led to a growing research area that employs LLMs as controllers to construct autonomous agents capable of effectively performing complex tasks Wang et al. (2024a). Building on these advances, recent state-of-the-art research has achieved remarkable success by designing LLM-based autonomous agents to automate the entire pipeline of literature review and generate long-form manuscripts Wang et al. (2024b); Liang et al. (2025).
Although these advances enable the automatic generation of human-like literature reviews with reduced manual effort, previous studies on multi-step models show that the inherent long-horizon workflow of such systems inevitably introduces the risk of compounding errors. In the workflow, small mistakes made at earlier steps can accumulate and propagate through subsequent steps, as errors that occur at a certain step affect the next state Wang et al. (2025a); Han and Zhang (2023); Venkatraman et al. (2015); Hejna et al. (2023). This paper investigates the phenomenon of error accumulation in automatic literature review generation. As a demonstration, we employ the large language model GPT-4.1 OpenAI (2025) to conduct a literature review generation workflow on the topic of EEG (electroencephalogram)-based emotion recognition, with a target length of 4,000 words. As shown in Fig. 1, when the retrieval stage fails to identify relevant literature, the resulting errors can mislead the outlining process. An irrelevant outline may further cause the drafted manuscript to deviate from the intended topic.
@@
However, directly applying these methods faces challenges in the context of literature review generation. Applying self-correction to the final output is difficult because errors have already accumulated across multiple stages, making the result hard to repair. Applying self-correction at each intermediate stage fixes errors within that stage but neglects context across stages. Thus, we propose to construct a memory mechanism that consolidates verified information along the long-horizon workflow using structured representations. Results of stages in the workflow are verified and consolidated into the memory, and subsequent steps expand from this foundation. Errors along the long horizon are filtered during consolidation and replaced during expanding before they can propagate, narrowing the search space around verified facts.
Based on these observations, we propose the Structure-Guided Memory Consolidation (SGMC) framework for automatic literature review generation. Specifically, Tree-Guided Memory ensures each outline level is grounded in actual literature, reducing topic drift. Hub-Guided Memory tightly aligns draft content with supporting evidence through iterative refinement. Self-Loop Memory introduces a dynamic checklist mechanism, enabling agents to learn from historical errors and proactively prevent error propagation.
Figure 2: The overall framework of Structure-Guided Memory Consolidation (SGMC) for literature review generation.
@@ 2 Related Work
Automating the summarization of research papers has long attracted interest. Early work used multi-document summarization techniques. In 2010, Hoang et al. introduced automatic related work summarization, extracting sentences from multiple articles to form an extractive summary Hoang and Kan (2010). Lu et al. constructed Multi-XScience, a dataset for generating related work sections based on a paper’s abstract and cited articles, which became a benchmark for later studies Lu et al. (2020). Chen et al. proposed a relation-aware multi-document encoder and a relation graph to capture cross-document dependencies for related work generation Chen et al. (2021). Xiao et al. developed PRIMERA, a pre-trained encoder-decoder model with objectives for linking and aggregating information across documents Xiao et al. (2022). Liu et al. proposed aligning sentences by category and using sparse transformers to organize information from references Liu et al. (2023).
Recent advances in LLMs have improved the scope, enabling end-to-end systems to generate literature review papers. Wang et al. introduced AutoSurvey, a framework with retrieval, outline generation, parallel subsection drafting, integration, and evaluation Wang et al. (2024b). Shao et al. developed STORM, which models the pre-writing stage by discovering diverse perspectives, simulating expert conversations, and curating information to create an outline Shao et al. (2024). Liang et al. designed SurveyX, an end-to-end solution for automated survey generation, covering online literature search, organization, and survey writing Liang et al. (2025). However, little attention has been paid to mitigating the risk of errors in generated outputs.
A response is post-hoc self-correction and refinement. Self-Refine Madaan et al. (2023) uses a single LLM as the generator, refiner, and feedback provider, iteratively improving outputs through self-generated feedback. Reflexion Shinn et al. (2023) reinforces language agents through linguistic feedback, where agents reflect on task feedback signals and maintain reflective text in an episodic memory buffer to induce better decision-making in subsequent trials. ReConcile Chen et al. (2024) enhances collaborative reasoning via multiple rounds of discussion among diverse LLM agents, employing a confidence-weighted voting mechanism to reach consensus. Chain-of-Note Yu et al. (2024) generates sequential reading notes for each retrieved document, enabling evaluation of their relevance before formulating the answer. RARR Gao et al. (2023) finds attribution for the output of any text generation model and post-edits the output to fix unsupported content while preserving the original output. However, in long-horizon literature review generation, accumulated errors may propagate too far for effective post-hoc repair, and directly applying these methods fails to provide reliable attribution and revision.
@@ 3 Methodology
As illustrated in Fig. 2, given a user instruction $U$, the proposed Structure-Guided Memory Consolidation (SGMC) framework orchestrates three memory structures to produce the final literature review output $Z$.
3.1 Tree-Guided Memory
The Tree-Guided Memory is designed for the stage of outlining the literature review and retrieving relevant references for each section of the outline. Without prior knowledge of the relevant references, the initial outline tends to be suboptimal. Conversely, without a determined direction, the literature retrieval process lacks focus. To mitigate compounding errors, we propose a hierarchical exploration strategy.
In detail, given a user instruction $U$, the agent $A_{M}$ based on a large language model is responsible for coordinating sub-agents using a Tree-Guided Memory. First, $A_{M}$ constructs a tree with the user instruction as the root node, representing the overall direction for exploring the field. This tree is expanded by identifying sub-directions of interest. For the root node, $A_{M}$ first invokes the agent $A_{S}$, which is responsible for constructing search queries and then calling search tools to retrieve references based on the user instruction, resulting in a set of relevant papers:
@@ 3.1 Tree-Guided Memory
where $L_{i}^{(0)}$ denotes the $i$-th piece of reference retrieved for the root node at tree depth $d=0$, and $I$ denotes the number of references retrieved at the root level.
Next, $A_{M}$ assigns the agent $A_{O}$, which is responsible for outlining, to determine the sub-outline based on the titles and abstracts of the retrieved literature:
$\{O_{1}^{(0)},O_{2}^{(0)},\ldots,O_{J}^{(0)}\}=A_{O}(\{L^{(0)}_{i}\}_{i=1}^{I})$ (2)
where $O_{j}^{(0)}$ denotes the $j$-th sub-outline at the root level, and $J$ is the total number of sub-outlines at tree depth $d=0$.
For each sub-outline node $O_{j}^{(d)}$ at depth $d$, $A_{M}$ decides whether further decomposition is needed according to the following criterion:
@@
$\{L^{(d+1)}_{1},L^{(d+1)}_{2},\dots,L^{(d+1)}_{I^{\prime}}\}$ $=A_{S}(O_{j}^{(d)})$ (4)
$\{O_{1}^{(d+1)},O_{2}^{(d+1)},\ldots,O_{J^{\prime}}^{(d+1)}\}$ $=A_{O}(\{L_{i}^{(d+1)}\}_{i=1}^{I^{\prime}})$
@@
where $I^{\prime}$ denotes the number of literature pieces retrieved at depth $d+1$, and $J^{\prime}$ denotes the number of sub-directions at depth $d+1$.
@@
This process constructs child nodes in the tree $\mathcal{T}$, where nodes $\{O_{1}^{(d+1)},O_{2}^{(d+1)},\ldots,O_{J^{\prime}}^{(d+1)}\}$ are the children of node $O_{j}^{(d)}$. The exploration process terminates when the following condition is met:
$\forall v\in\mathcal{V}:\phi(v)=0\quad\text{or}\quad d(v)=d_{\max}$ (5)
@@
where $\mathcal{V}$ denotes all leaf nodes in the outline tree, $d(v)$ denotes the depth of node $v$ in the tree $\mathcal{T}$, and $d_{\max}$ is the maximum allowed depth. The resulting tree $\mathcal{T}$ provides a hierarchical outline, with associated literature at each level, for downstream generation.
3.2 Hub-Guided Memory
@@
The Hub-Guided Memory is designed to extract relevant snippets from the literature and draft the review manuscript. Without clear target content, it is difficult to locate the necessary evidence in the literature. Conversely, without sufficient supporting evidence, the manuscript may lack grounding. To mitigate compounding errors, we propose an iterative exploitation strategy.
@@
In detail, given the outline tree $\mathcal{T}$, $A_{M}$ initiates the Hub-Guided Memory. For each leaf node in the tree, $A_{M}$ identifies its parent outline node, denoted as $\mathcal{O}_{k}$, and the associated reference set $\{\mathcal{L}_{i}\}_{i=1}^{I_{k}}$, where $I_{k}$ is the number of references associated with outline node $\mathcal{O}_{k}$, connected to the hub node. Then, $A_{M}$ coordinates sub-agents to aggregate information from references to the hub node through iterative cycles. At the initial iteration $t=0$, the agent $A_{L}$, which is responsible for locating, extracts relevant information based on the outline node $\mathcal{O}_{k}$ and the literature set $\{\mathcal{L}_{i}\}_{i=1}^{I_{k}}$:
$\{F_{1}^{(0)},F_{2}^{(0)},\ldots,F_{N}^{(0)}\}=A_{L}(\{\mathcal{L}_{i}\}_{i=1}^{I_{k}},\mathcal{O}_{k})$ (6)
@@
where $F_{n}^{(0)}$ denotes the $n$-th note snippet extracted from the reference set $\{\mathcal{L}_{i}\}_{i=1}^{I_{k}}$ relevant to the outline node $\mathcal{O}_{k}$, and $N$ is the total number of snippets.
@@
The agent $A_{D}$, which is responsible for drafting, composes the initial draft as follows:
$M^{(0)}_{k}=A_{D}(\{F^{(0)}_{n}\}_{n=1}^{N},\mathcal{O}_{k})$ (7)
@@
where $M^{(0)}_{k}$ represents the initial manuscript for the outline node $\mathcal{O}_{k}$, incorporating the content of all its child outline nodes.
@@
For subsequent iterations $t>0$, $A_{L}$ re-examines the literature to extract new snippets to update the hub and improve the current draft:
$\{F_{1}^{(t)},F_{2}^{(t)},\ldots,F_{N^{\prime}}^{(t)}\}=A_{L}(\{\mathcal{L}_{i}\}_{i=1}^{I_{k}},M^{(t-1)}_{k})$ (8)
@@
where $M^{(t-1)}_{k}$ denotes the draft content from the previous iteration, $\{F^{(t)}_{n}\}_{n=1}^{N^{\prime}}$ is the set of newly extracted note snippets, and $N^{\prime}$ is the number of snippets at iteration $t$. $A_{D}$ refines the section draft by incorporating the new evidence and the previous draft:
$M^{(t)}_{k}=A_{D}(\{F^{(t)}_{n}\}_{n=1}^{N^{\prime}},\mathcal{O}_{k},M^{(t-1)}_{k})$ (9)
where $M^{(t)}_{k}$ is the updated content, and $M^{(t-1)}_{k}$ serves as the previous context for refinement.
@@
The iterative process for each outline node $\mathcal{O}_{k}$ terminates when the following condition is met:
$\psi(M^{(t)}_{k},M^{(t-1)}_{k})>\epsilon\quad\text{or}\quad t=t_{\max}$ (10)
@@
where $\psi(\cdot,\cdot)$ denotes the similarity between consecutive iterations of the draft, measured by the ROUGE-1 score Lin (2004), $\epsilon$ is a similarity threshold that indicates no further improvements can be made, and $t_{\max}$ is the maximum number of allowed iterations.
@@
The final content $M^{(t^{*})}_{k}$, where $t^{*}$ denotes the stopping iteration for outline node $\mathcal{O}_{k}$, serves as the final draft. Finally, $A_{M}$ aggregates the finalized drafts for all sections and delivers the completed literature review to the user.
3.3 Self-Loop Memory